# Robust brain age estimation from structural MRI with contrastive learning

### v1

@@ 4.2 Experimental setup
Across all of the experiments we employ a 5-fold CV scheme, at the patient level, where the splits are stratified by age w.r.t to the original distribution. In experiments targeting only brain age prediction, we consider only HC patients (the distribution is presented in Tab. 3), whereas in experiments targeting cognitive decline we consider all patients (HC, AD, and MCI). As network architecture, we employ the 3D implementations of ResNet-18 with 33.2M parameters. We stick to ResNet-18 as in our previous work (Barbano et al., 2023a), we show that bigger models do not provide better results. In line with Barbano et al. (2023a), we use the Adam optimizer, with an initial learning rate of $10^{-4}$ decayed by a factor of $0.9$ every 10 epochs, and a weight decay of $5\times 10^{-5}$. We use a batch size of 32, and train for a total of 300 epochs. Our experiments are implemented in PyTorch and run on the Leonardo and Jean Zay clusters, using A100 and V100 NVIDIA GPUs.
@@
For every model, we evaluate the mean absolute error (MAE) and the $R^{2}$ coefficient on the test set of each dataset. Furthermore, the OpenBHB challenge computes the MAE on both the internal (Int. MAE) and external (Ext. MAE) test sets and the balanced accuracy (BAcc) on site prediction, training a logistic regression on the model representations. The final challenge score is then computed as $\mathcal{L}_{c}=\text{BAcc}^{0.3}\cdot\text{MAE}_{\text{ext}}$. On AD and MCI patients, we also consider the ROC curves computed on BAG, and fine-tuning accuracy for classification.
@@ 4.3 Proposed method
In Contrastive Learning (CL), given an ensemble of samples (here images) $x\in\mathcal{X}$, one looks for a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that maps similar (positive) samples $x^{+}$ close in the representation space (a $\mathbb{S}^{d-1}$ hypersphere) while repelling negative ones $x^{-}$. Supervised CL, e.g., SupCon (Khosla et al., 2020) or $\epsilon$-SupInfoNCE (Barbano et al., 2023b), operates with discrete labels $y$ by defining positive pairs as samples from the same class and negatives otherwise. Given an anchor sample $x_{i}$ with latent representation $f(x_{i})$, supervised CL aligns it with all samples of the same class and repels it from those of different classes. While this formulation works well for classification, it is not naturally suited to regression, where labels are continuous (i.e., $y\in\mathbb{R}$) and no clear boundary separates positives from negatives samples. To solve this issue, in Barbano et al. (2023a) we propose to use a degree of similarity defined as a kernel function $K$ on the continuous labels $y$:
$w_{k}=K(y-y_{k}),\quad\text{where }0\leq w_{k}\leq 1.$ (1)
Here, $w_{k}$ encodes how similar the sample $x_{k}$ is to the anchor $x_{i}$ in label space. As for $K$, we choose a Radial Basis Function (RBF) parametrized by $\sigma$, chosen by the user, but other choices can be made (Barbano et al., 2023a). Our goal becomes to learn a mapping $f$ such that samples with high degree of similarity $w_{k}$ are embedded close to the anchor in the representation space $\mathbb{S}^{d-1}$, while those with low $w_{k}$ are pushed farther apart.
@@
We begin from the general contrastive condition (Barbano et al., 2023b):
@@
$\mathop{\mathrm{arg\,min}}_{f}\sum_{k}\max(0,\{s^{-}_{t}-s^{+}_{k}\}_{\forall t})=-\sum_{k}\log\left(\frac{\exp s^{+}_{k}}{\exp s^{+}_{k}+\sum_{t}\exp s^{-}_{t}}\right)$ (3)
This formulation is well suited for classification but not for regression, since one would need well separated classes (thus discrete labels). In Barbano et al. (2023a), we propose different formulations of the starting condition (2) that are more suited for regression (continuous labels). We describe them here.
$\mathcal{L}^{y\text{-aware}}$ (Weakly-supervised contrastive loss)
A straightforward idea to tackle regression with continuous labels is to softly weight samples by their similarity to the anchor $x_{i}$:
$\frac{w_{k}}{\sum_{j}w_{j}}(s_{t}-s_{k})\leq 0\quad\forall j,k,t\neq k\in A(i),$ (4)
where $A(i)$ is the set of all samples in the batch excluding $i$, and the weights are normalized to sum to one. This condition can be transformed into an optimization problem using LogSumExp:
$\mathcal{L}^{y\text{-aware}}$ $=-\sum_{k}\frac{w_{k}}{\sum_{t}w_{t}}\log\left(\frac{\exp(s_{k})}{\sum_{t=1}^{N}\exp(s_{t})}\right).$ (5)
This formulation, originally proposed in Dufumier et al. (2021), generalizes contrastive learning to continuous attributes. However, its denominator acts as a uniformity term (Wang & Isola, 2020) that may cause the model to repel samples that are more similar to the anchor than $x_{k}$, which is undesirable in regression. This is because the summation over $t$ includes all samples in the minibatch, independently of their similarity to the anchor.
$\mathcal{L}^{thr}$ (Threshold-aware loss)
To address this, we propose to include in the denominator only those samples that are less similar to the anchor than $x_{k}$ (i.e., with $w_{t}<w_{k}$):
$\mathcal{L}^{thr}=-\sum_{k}\frac{w_{k}}{\sum_{t}\delta_{w_{t}<w_{k}}w_{t}}\log\left(\frac{\exp(s_{k})}{\sum_{t\neq k}\delta_{w_{t}<w_{k}}\exp(s_{t})}\right).$ (6)
This prevents repelling samples more similar than $x_{k}$, but still, due to the presence of the exponential, it emphasizes the closest negative just under the threshold, which is not ideal.
$\mathcal{L}^{exp}$ (Exponential contrastive loss)
To overcome this limitation, we introduce an exponential formulation that modulates repulsion strength based on dissimilarity:
$\mathcal{L}^{exp}=-\frac{1}{\sum_{t}w_{t}}\sum_{k}w_{k}\log\left(\frac{\exp(s_{k})}{\sum_{t\neq k}\exp(s_{t}(1-w_{t}))}\right).$ (7)
@@
5 Related Works
Contrastive learning
When label information is not available, instance-based discrimination approaches (Chen et al., 2020) have gained considerable success, almost closing the gap with its supervised counterpart. Most recent research in CL has focused on designing the right data augmentations (Chen et al., 2020; Zhang & Ma, 2022), contrastive losses (Chen et al., 2020; Caron et al., 2021; Zbontar et al., 2021; Grill et al., 2020), memory banks (He et al., 2020; Dwibedi et al., 2021), prototypes (or clusters) (Caron et al., 2020; Li et al., 2020) typically for downstream classification, detection or segmentation tasks, leaving open the question for regression tasks. When weak labels are available (either discrete or continuous), kernel-based approaches have been explored for CL (Dufumier et al., 2021, 2023), using a re-weighting scheme to adapt the alignment term between samples with close weak labels. When continuous labels are available, ranking-based approaches (Zha et al., 2023) have recently gained traction by showing strong generalization results on regression tasks.
Brain age estimation
Estimating the chronological age from brain imaging in healthy individuals has been a challenge for machine learning models for the past 10 years. The first traditional models (linear, SVM) were trained on handcrafted features extracted from structural MRI (cortical thickness, gray matter volume, surface area) (Cole et al., 2015; Franke & Gaser, 2012). Then, more complex deep learning models, such as CNNs (Cole et al., 2017) or Transformers (He et al., 2021), have been deployed to automatically extract brain features from raw imaging. These brain age models were used to derive a brain age gap correlated to common brain disorders such as schizophrenia, bipolar disorder, ASD, and ADHD (Kaufmann et al., 2019) and heritable (Cole et al., 2017; Jónsson et al., 2019). Nevertheless, recent studies (Dufumier et al., 2024) showed that DL models generalize poorly to new MRI sequences/acquisitions, and they still underperform on populations with different demographic characteristics.

### v2

@@ 3.2 Experimental setup
Across all of the experiments we employ a 5-fold CV scheme, at the subject level, where the splits are stratified by age w.r.t to the original distribution. In experiments targeting only brain age prediction, we consider only HC patients (the distribution is presented in Tab. 2), whereas in experiments targeting cognitive decline we consider all patients (HC, AD, and MCI). As network architecture, we employ the 3D implementations of ResNet-18 with 33.2M parameters, as previous works (Dufumier et al., 2021a, 2022, 2024; Barbano et al., 2023a, 2026; Siegel et al., 2025; Yang et al., 2025) show that different model sizes and architectures do not provide better results. In addition to our proposed losses, we also perform experiments with RNC (Zha et al., 2023) and AdapN (Träuble et al., 2024b). For all methods, in line with previous works (Barbano et al., 2023a; Träuble et al., 2024b), we use the Adam optimizer, with an initial learning rate of $10^{-4}$ decayed by a factor of $0.9$ every 10 epochs, and a weight decay of $5\times 10^{-5}$. We use a batch size of 32, and train for a total of 300 epochs. For contrastive losses we employ a temperature of 2.0 for RNC (Zha et al., 2023) and 0.1 for all others. Our experiments are implemented in PyTorch and run on the Leonardo and Jean Zay clusters on A100 and V100 NVIDIA GPUs.
@@
We evaluate the mean absolute error (MAE), the $R^{2}$ coefficient pearson correlation $r$ on the test set of each dataset. Furthermore, the OpenBHB challenge computes the MAE on both the internal (Int. MAE) and external (Ext. MAE) test sets and the balanced accuracy (BAcc) on site prediction, training a logistic regression on the model representations. The final challenge score is then computed as $\mathcal{L}_{c}=\text{BAcc}^{0.3}\cdot\text{MAE}_{\text{ext}}$. On AD and MCI patients, we also consider the ROC curves computed on BAG, and fine-tuning accuracy for classification.
Table 3: Brain age regression results. Single means that the models were trained and tested on each dataset separately; Merge means that the models were trained on the combined healthy samples. OpenBHB Challenge Results report balanced accuracy (BAcc), internal MAE, external MAE, and the composite score $\mathcal{L}_{c}$. For single training, the challenge evaluation was performed on models trained on OpenBHB.
OpenBHB ADNI ICBM OASIS3 OpenBHB Challenge Results
Method Train MAE $R^{2}$ r (corr.) MAE $R^{2}$ r (corr.) MAE $R^{2}$ r (corr.) MAE $R^{2}$ r (corr.) BAcc Int. MAE Ext. MAE $\mathcal{L}_{c}\,(\downarrow)$
L1 Single 2.73 0.92 0.96 3.64 0.58 0.78 4.33 0.75 0.87 4.32 0.65 0.82 6.7 2.67 4.18 1.82
$\mathcal{L}^{y-aware}$ (Dufumier et al., 2021c) 2.65 0.92 0.96 3.47 0.61 0.79 4.21 0.76 0.87 4.26 0.65 0.82 6.60 2.66 4.10 1.82
$\mathcal{L}^{threshold}$ (Barbano et al., 2023a) 2.66 0.93 0.96 3.36 0.63 0.65 5.27 0.57 0.86 4.18 0.66 0.82 5.73 2.95 4.10 1.74
$\mathcal{L}^{exp}$ (Barbano et al., 2023a) 2.66 0.93 0.96 3.44 0.62 0.79 4.11 0.77 0.88 4.22 0.66 0.82 5.1 2.55 3.76 1.54
RNC (Zha et al., 2023) 3.78 0.86 0.56 3.73 0.56 0.45 4.57 0.72 0.51 4.48 0.63 0.49 4.9 3.63 4.30 1.74
$\mathcal{L}^{AdapN}$ (Träuble et al., 2024b) 2.54 0.93 0.58 3.47 0.61 0.79 4.59 0.71 0.86 4.30 0.65 0.82 5.53 2.61 4.16 1.75
L1 Merge 2.59 0.93 0.97 3.45 0.61 0.79 3.77 0.82 0.91 4.08 0.68 0.84 9.33 2.42 2.43 1.19
$\mathcal{L}^{y-aware}$ (Dufumier et al., 2021c) 2.70 0.91 0.96 3.40 0.63 0.80 3.73 0.82 0.92 4.08 0.67 0.84 7.3 2.61 2.53 1.15
$\mathcal{L}^{threshold}$ (Barbano et al., 2023a) 3.11 0.85 0.94 3.43 0.62 0.80 3.83 0.80 0.91 4.11 0.66 0.84 6.0 2.90 2.21 0.95
$\mathcal{L}^{exp}$ (Barbano et al., 2023a) 2.77 0.91 0.95 3.39 0.63 0.80 3.84 0.81 0.91 4.05 0.68 0.85 5.40 2.71 2.25 0.93
RNC (Zha et al., 2023) 4.92 0.77 0.70 5.57 -0.01 0.42 4.95 0.70 0.67 6.39 0.19 0.38 2.7 5.08 6.13 2.07
$\mathcal{L}^{AdapN}$ (Träuble et al., 2024b) 2.71 0.92 0.96 3.42 0.62 0.80 3.71 0.83 0.92 3.96 0.69 0.84 6.37 2.54 2.36 1.03
@@ 3.3 Proposed method
In this section, we present our contrastive learning framework for regression, originally introduced in Barbano et al. (2023a), leading to the formulation of our proposed loss function $\mathcal{L}^{exp}$.
@@
In contrastive learning, given a collection of samples (here, images) $x\in\mathcal{X}$, the objective is to learn a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that projects similar, or positive, samples $x^{+}$ close together in a shared embedding space (the hypersphere $\mathbb{S}^{d-1}$), while pushing negative samples $x^{-}$ apart (Khosla et al., 2020; Barbano et al., 2023b). For a given anchor sample $x_{i}$ with latent representation $f(x_{i})$, the model encourages high similarity with embeddings of samples sharing the same label and low similarity with those from other categories. This formulation has proven highly effective for classification problems; however, it is not directly applicable to regression, where labels are continuous (i.e., $y\in\mathbb{R}$) and the distinction between positive and negative pairs is inherently ambiguous. To solve this issue, in Barbano et al. (2023a) we propose to use a degree of similarity defined as a kernel function $K$ on the continuous labels $y$:
$w_{k}=K(y-y_{k}),\quad\text{where }0\leq w_{k}\leq 1$ (1)
Here, $w_{k}$ encodes how similar the sample $x_{k}$ is to the anchor $x_{i}$ in label space. As for $K$, we choose a Radial Basis Function (RBF) parametrized by $\sigma$, chosen by the user, but other choices can be made. Our goal becomes to learn a mapping $f$ such that samples with high degree of similarity $w_{k}$ are embedded close to the anchor in the representation space $\mathbb{S}^{d-1}$, while those with low $w_{k}$ are pushed farther apart.
@@
From this idea, we define different loss functions. We begin from the general contrastive condition (Barbano et al., 2023b):
@@
$\mathop{\mathrm{arg\,min}}_{f}\sum_{k}\max(0,\{s^{-}_{t}-s^{+}_{k}\}_{\forall t})\approx-\sum_{k}\log\left(\frac{\exp s^{+}_{k}}{\exp s^{+}_{k}+\sum_{t}\exp s^{-}_{t}}\right)$ (3)
This formulation is well suited for classification but not for regression, since one would need well separated classes (thus discrete labels).
$\mathcal{L}^{y\text{-aware}}$ (Dufumier et al., 2021c)
A straightforward idea to tackle regression with continuous labels is to softly weight alignment of samples by their similarity to the anchor:
$\mathcal{L}^{y\text{-aware}}$ $=-\sum_{k}\frac{w_{k}}{\sum_{t}w_{t}}\log\left(\frac{\exp(s_{k})}{\sum_{t=1}^{N}\exp(s_{t})}\right)$ (4)
This loss generalizes contrastive learning to continuous attributes. However, its denominator uniformly repels samples apart from the anchor, disregarding their similarity, as measured by the weights $(w_{t})_{t\neq k}$. This might be suboptimal for regression.
Non-uniform hard thresholding
To address this, we propose to include in the denominator only samples with lower degrees of similarity than $x_{k}$, i.e., with $w_{t}<w_{k}$:
$\mathcal{L}^{thresold}=-\sum_{k}\frac{w_{k}}{\sum_{t}\delta_{w_{t}<w_{k}}w_{t}}\log\left(\frac{\exp(s_{k})}{\sum_{t\neq k}\delta_{w_{t}<w_{k}}\exp(s_{t})}\right)$ (5)
This prevents repelling samples more similar than $x_{k}$. Nonetheless, because of the exponential in the denominator, it sill emphasizes the closest negative just under the threshold, which is not ideal.
Non-uniform soft thresholding
In Barbano et al. (2023a), we introduce an exponential formulation that modulates repulsion strength based on dissimilarity:
$\mathcal{L}^{exp}=-\frac{1}{\sum_{t}w_{t}}\sum_{k}w_{k}\log\left(\frac{\exp(s_{k})}{\sum_{t\neq k}\exp(s_{t}(1-w_{t}))}\right)$ (6)
@@
4 Results
(a) External MAE vs training size (scaling)
(b) Site bias vs training size. $\mathcal{L}^{exp}$ stays consistent.
Fig. 1: Scaling contrastive brain-age models with data size. (a) External mean absolute error (MAE) decreases as training size increases; all methods besides RnC benefit from more data, reaching a MAE between 2.21 and 2.53. (b) Balanced accuracy (B. Acc) in predicting scanner site increases for most methods as training size grows, indicating increased site bias. In contrast, $\mathcal{L}^{exp}$ remains stable and low across all sizes, demonstrating robustness to site-specific confounds. On the other hand, the decrease in B. Acc for RnC can be attributed to the degraded performance in brain age prediction when datasets are merged (e.g., the model collapsed and did not learn meaningful representations). In fact RnC shows a trend similar to other methods when trained on OpenBHB only. We hypothesize this is due to the unweighted alignment term and to the high temperature value used by RnC, which could focus the repulsion on samples coming from different dstributions (e.g., ADNI and OASIS subjects are older), leading to a latent space which does not capture well biological and age variability.
(a) Brain Age Gap (BAG) distribution for HC, MCI, and AD classes across all datasets. The shaded areas represent the age delta distribution (histogram) for each class. The horizontal black line in the middle highlights the mean value (explicitly annotated for readability).
(b) ROC plots for using BAG as a proxy for detecting AD patients.
Fig. 2: Using BAG as a proxy for detecting neurodegeneration. (a) BAG distribution for HC, stable MCI (sMCI), progressive MCI (pMCI), and AD patients across all datasets. All methods consistently show increased BAG in pMCI and AD groups, indicating accelerated brain aging. (b) ROC curves evaluating the ability of BAG to separate HC and AD subjects in ADNI and OASIS datasets. Contrastive methods ($\mathcal{L}^{exp},\mathcal{L}^{threshold},\mathcal{L}^{yaware}$) match or slightly outperform the L1 baseline. (c) Longitudinal BAG trajectories (with $\mathcal{L}^{exp})$ on ADNI, showing that pMCI and AD subjects exhibit increased BAG over time, while HC and sMCI groups remain stable. This confirms that contrastive models capture disease progression dynamics.
4.1 Scaling training size improves generalization
In Tab. 3 we report results of brain age regression across the different datasets employed. We experiment with two pre-training schemes, namely, we first train and test on the same dataset (single), and then we train on the combined healthy population across all datasets and test on each test set separately (merge). As we can note from the results, the overall error decreases when pretraining on the larger dataset. We also benchmark our models on the public OpenBHB challenge, where we significantly improve the results in terms of generalization (external MAE). With $\mathcal{L}^{exp}$, we achieve state-of-the-art results in the challenge, scoring an impressive 0.93, compared to the previous SOTA of 1.54. By increasing the training set size, we are able to almost halve the external prediction error, meaning that the models generalize better. Fig. 1(a) shows models scaling in terms of external MAE compared to the training set size. As we can see, almost all methods benefit from increasing the pre-training size. This is consistent with related literature (Schulz et al., 2024), showing that deep models follow exponential laws in brain age prediction from T1 structural MRIs. This does always hold for RnC, which collapses when datasets are merged. We hypothesize this is due to the unweighted alignment term and to the high temperature value (Zha et al., 2023).
4.2 $\mathcal{L}^{exp}$ is more invariant to site bias
By closely inspecting results in Tab. 3, we notice an interesting phenomenon: all models pre-trained on the merged dataset, other than $\mathcal{L}^{exp}$, exhibit a higher balanced accuracy (Bacc) in retrieving the acquisition site of the scan, meaning that as the training size gets larger, they become more biased towards the site. $\mathcal{L}^{exp}$ is the only method which preserve a final balanced accuracy below 6%. To illustrate these results in more detail, in Fig. 1(b), we show how the final balanced accuracy scales with respect to the size of the training data, in terms of acquisition sites and datasets. As we can observe from the results, our proposed loss $\mathcal{L}^{exp}$ shows a balanced accuracy almost constant at different training size, whereas most other methods are heavily influenced by the training composition, and tend to become more biased as diversity in the data increases. We hypothesize that this is due to the difference in the formulation of $\mathcal{L}^{exp}$ with respect to the other losses, such as y-aware. The presence of the exponential in the uniformity term of y-aware (Dufumier et al., 2021c) may force some samples to be pushed apart from the anchor more than they should. This could result in a less “compact” latent space, that might capture more variance in the input data, such as noise from the site effect. On the other hand, $\mathcal{L}^{exp}$ may avoid this phenomenon as the samples are repelled according to the kernel value (Barbano et al., 2023a), achieving latent representations that are more invariant to such noise.
Table 4: Downstream results (AD detection) and cross-dataset generalization in terms of balanced accuracy (%). Models are pre-trained on OpenBHB, finetuned on ADNI or OASIS, and tested on each dataset. Results show that CL methods are always competitive or better than L1 pre-training.
ADNI $\rightarrow$ ADNI ADNI $\rightarrow$ OASIS OASIS $\rightarrow$ ADNI OASIS $\rightarrow$ OASIS Avg.
L1 86.8% $\pm$2.8% 50.0% $\pm$0.0% 53.5% $\pm$11.3% 62.0% $\pm$4.4% 63.1%
$\mathcal{L}^{yAware}$ 85.0% $\pm$2.5% 60.7% $\pm$7.2% 71.1% $\pm$13.5% 67.0% $\pm$5.8% 70.9%
$\mathcal{L}^{threshold}$ 84.9% $\pm$2.4% 51.9% $\pm$1.9% 65.1% $\pm$8.9% 67.8% $\pm$7.8% 67.4%
$\mathcal{L}^{expw}$ 85.7% $\pm$3.0% 68.8% $\pm$6.7% 66.1% $\pm$12.5% 71.8% $\pm$5.0% 73.1%
$\mathcal{L}^{AdapN}$ 85.4% $\pm$2.4% 51.4% $\pm$1.6% 82.8% $\pm$2.4% 62.7% $\pm$4.9% 70.6%
RnC 85.8% $\pm$2.6% 55.1% $\pm$4.1% 83.5% $\pm$1.6% 64.2% $\pm$5.1% 72.1%
Fig. 3: Calibration curves show that all fine-tuned models are reasonably calibrated on the downstream task, with an Average Calibration Error (ECE) slightly higher than 10%.
Fig. 4: Correlation between downstream accuracy (ADNI) and brain age prediction error. Contrastive methods such as $\mathcal{L}^{exp}$ benefit from lower MAE.
4.3 CL is a reliable estimator of accelerated aging
Next, we aim to assess whether CL is a reliable estimator to detect accelerated aging in subjects affected by cognitive impairment and neurodegenerative diseases such as Alzheimer’s Disease. Brain age is widely recognized as a biomarker of different brain disorders (Kaufmann et al., 2019), and many works proposed deep learning pipelines for brain age estimation (Cole et al., 2017; He et al., 2021). These models are typically trained with a supervised loss such as L1 and have shown robust performance in highlighting accelerated aging in affected patients. To assess whether our proposed CL framework also exhibits this property, we employ data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) with 1754 subjects and from OASIS-3 with 685 subjects for a total of 2439 subjects. The analysis is presented in Fig. 2. First, in Fig. 2(a), we show the distribution of brain age gap (BAG) - the difference between the model’s predicted age and the subject’s chronological age - across all groups HC, stable mild cognitive impairment (sMCI), progressive mild cognitive impairment (pMCI), and AD. The obtained results are mostly consistent among all methods, showing that contrastive losses are comparable to standard L1 supervised models. Furthermore, we notice that L1 exhibits a higher spread on the pMCI and AD classes, compared to CL methods. In Fig. 2(b) we report the ROC curves for detecting AD patients on ADNI and OASIS. Once more, the results show that contrastive learning methods are comparable to standard L1 supervised baselines, even with a slight edge (AUC of 0.82 vs 0.81 on ADNI and 0.73 vs 0.72 on OASIS). Lastly, we assess the longitudinal reliability of $\mathcal{L}^{exp}$ by calculating the change in BAG during follow-ups for ADNI subjects. As expected from Franke & Gaser (2012), we observe an increase of BAG for AD and pMCI patients, meaning an increased brain age during follow-ups; whereas HC and sMCI subjects remain stable.
4.4 Do better brain age models translate to better diagnostic models?
Finally, we also analyze whether more accurate brain age prediction models can translate to better diagnostic models. To do so, we finetune the models pre-trained on OpenBHB using ADNI and OASIS data. Results are reported in Tab. 4. Results show that CL is almost always better than L1 pre-training. The proposed $\mathcal{L}^{expw}$ achieves the best generalization in two out of four transfer directions, reaching accuracies of 68.8% and 71.8%, respectively, and achieves the highes average accuracy. Similarly, RnC and $\mathcal{L}^{AdapN}$ perform competitively when transferring from OASIS to ADNI, with accuracies exceeding 82%. On the other hand, L1 achieves results closer to a random guess in both cross-dataset generalization tasks. These results confirm that contrastive pre-training yields more transferable representations that maintain discriminative power across acquisition sites and cohorts. In terms of model calibration (Fig. 3), all methods remain reasonably well-calibrated, with Expected Calibration Error (ECE) values around 10–12% on ADNI. We also report the downstream accuracy in classifying AD and HC versus the original brain age MAE in Fig. 4. Consistently with related literature (Bashyam et al., 2020), we find that standard L1 supervised models do not really benefit on average from lower age prediction error, showing a correlation between age MAE and classification accuracy of $r=0.28$. In fact, Bashyam et al. (2020); Schulz et al. (2025) note that tighter fitting models on age naturally produce estimates that offer little information beyond age, and moderately fitting brain age models obtain significantly higher differentiation in a transfer-learning setup. On the contrary, our results seem to suggest that models pre-trained using CL can learn complex and meaningful representations that can benefit downstream tasks. Specifically, $\mathcal{L}^{exp}$ shows an almost total correlation of $r=0.92$ between age prediction error and downstream accuracy, and almost all CL methods exhibit a significantly higher slope than the L1 baseline (meaning that the increase in accuracy is more rapid). This shows that CL is less prone to overfitting the brain age prediction task. This is an important finding, as it may lay the foundation for robust pre-trained pipelines on neuroimaging data.
5 Discussion
On all benchmarks, training on the merged healthy cohort reduced the error relative to the single-dataset baseline (Tab. 3). The gains are most pronounced on the OpenBHB challenge, where the external MAE nearly halves on average. This mirrors the scaling laws recently reported for L1-supervised models (Schulz et al., 2024), and shows that CL obeys similar dynamics. Importantly, Fig. 1(a) shows that performance does not saturate in the range we explored, suggesting further room for improvement as larger neuroimaging datasets become available.
$\mathcal{L}^{exp}$ mitigates site bias at scale
While every loss benefited from more data, only $\mathcal{L}^{exp}$ maintained a low balanced accuracy ($6\%$) in predicting acquisition site after pre-training (Tab. 3, Fig. 1(b)). We hypothesize that the weighted repulsion term in $\mathcal{L}^{exp}$ prevents over-dispersion of negatives, yielding a latent space that captures biologically relevant variance while suppressing site-specific factors. By contrast, the fixed uniformity term in $y$-Aware and $\mathcal{L}^{threshold}$ can force hard negatives to drift apart excessively, leaving room for confounding factors to be learned. In this regard, L1 shows the strongest bias towards site. In practice, this robustness is crucial when models are deployed across centres that differ in hardware and protocol.
CL pre-training is a reliable detector of accelerated ageing
All CL models replicated the well-established pattern of higher brain-age gaps in pMCI and AD groups (Fig. 2(a)), also showing slightly lower variance than the L1 baseline. BAG-based ROC curves (Fig. 2(b)) equal or surpass the L1 reference (AUC 0.82 vs. 0.81 on ADNI; 0.73 vs. 0.72 on OASIS-3). Moreover, longitudinal trajectories on ADNI confirm that $\mathcal{L}^{exp}$ captures progressive neurodegeneration (Fig. 2(c)). Taken together, these results indicate that CL embeddings preserve clinically meaningful signals rather than merely fitting chronological age.
For L1 models, tighter age fits do not necesserily predict better HC/AD separation, consistent with prior reports that over-fitting to age harms downstream transfer (Bashyam et al., 2020; Tan et al., 2025). Interestingly, the opposite seems to hold for $\mathcal{L}^{exp}$ ($r=0.92$). We argue that CL pre-training encourages more general representations, whereas strongly supervised objectives push all capacity toward age alone. This finding positions CL models as promising foundation models for a broad range of neuroimaging tasks.
6 Conclusions
In this work, we performed a thorough assessment of state-of-the-art contrastive learning methods for regression on structural MRI data. Employing a dataset of over 20,000 scans, we show that contrastive learning improves both regression error and transferability compared to standard L1-supervised deep learning models. Scaling contrastive pre-training with the proposed $\mathcal{L}^{exp}$ loss establishes new state-of-the-art performance on brain-age prediction and, crucially, yields embeddings that (i) are robust to scanner variation, (ii) capture accelerated ageing in neurodegenerative disease, and (iii) transfer monotonically to clinical classification tasks. These properties make CL a strong candidate for building large neuroimaging foundation models that can underpin future precision-medicine pipelines.
Limitations
First, although we aggregate over 20k scans and OpenBHB gathers data from different populations, racial and ethnic diversity remains limited, especially in the downstream tasks; future work should verify that the invariance of $\mathcal{L}^{exp}$ extends also to demographic shifts. Second, our analysis focuses on T1-weighted MRI; multi-modal extensions (e.g., diffusion or functional MRI) could further enrich the learned representations. Third, we treated the site-prediction task as a proxy for scanner bias; a more comprehensive assessment might include vendor, field strength, and pulse-sequence metadata.